OpenStudy (anonymous):

HELP !! PLEASE!! Solve Log(little 8)2=x without calculator PLEASE

OpenStudy (anonymous):

\[\log_{8} 2=x\]

OpenStudy (jhannybean):

\[\large \log_a (n) = y \implies a^y = x\]\[a=8~,~ 2=n~,~ x=y\]

OpenStudy (jhannybean):

So you'll have \[\large 8^x = 2\]\[\large 8= 2^3\]\[\large (2^3)^x=2\]\[\large 2^{3x}=2\]Take \(\log_2\) of each side to get rid of the base so you can solve for the exponents.\[\log_2(2^{3x})= \log_2(2)\]Remember, \(\log_a (a) = 1\)

OpenStudy (anonymous):

okay so it gives you \[\log4^{3x}=\log4\]

OpenStudy (usukidoll):

are we just solving for x like find the value that makes \[\log_82=x\] true?

OpenStudy (usukidoll):

that's 3. 2^x = 8

OpenStudy (usukidoll):

so what x value produces 8 on the right? Well if I take the base 2 to the 3rd power I will have 2 x 2 x 2 = 8 or 2^3=8/ That's about it.

OpenStudy (anonymous):

Also: \[ \log_a(b) = \frac{\log_b(b)}{\log_b(a)} = \frac{1}{\log_b(a)} \]Interesting.

OpenStudy (anonymous):

So \[ \log_8(2) = \frac1{\log_2(8)} = \frac1{\log_2(2^3)} =\frac1{3} \]

OpenStudy (jhannybean):

Smarties.

OpenStudy (anonymous):

so the answer is 1/3 ? sorry for late reply